# Generalized Linear Bandits with Limited Adaptivity

### v4

@@ A.1 Additional Notation
$\gamma(\lambda)=24RS\left(\sqrt{\log{\left(T\right)}+d}+\frac{R\left(\log{\left(T\right)}+d\right)}{\sqrt{\lambda}}\right)+2S\sqrt{\lambda}.$ (9)
@@ A.2 Concentration Inequalities and Confidence Intervals
$\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{{\mathbf{H}}^{*}}}\leq 24RS\left(\sqrt{\log{\left(T\right)}+d}+\frac{R\left(\log{\left(T\right)}+d\right)}{\sqrt{\lambda}}\right)+2S\sqrt{\lambda}$ (11)
@@
[…]
${\mathbf{G}}$ $\coloneqq\sum_{j=1}^{s}\alpha(x,\theta^{*},\widehat{\theta})x_{j}x_{j}^{\mathsf{T}}+\frac{\lambda}{1+2RS}{\mathbf{I}}$
@@
$\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{{\mathbf{H}}^{*}}}$ $\leq\sqrt{(1+2RS)}\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{\mathbf{G}}}$
@@
$=\sqrt{1+2RS}\left\lVert\sum_{j=1}^{s}\left(\langle\theta^{*},x_{j}\rangle-\langle\widehat{\theta},x_{j}\rangle\right)\alpha(x,\theta^{*},\widehat{\theta})x_{j}+\frac{\lambda}{1+2RS}\left(\theta^{*}-\widehat{\theta}\right)\right\rVert_{{\mathbf{G}}^{-1}}$
$\leq\sqrt{1+2RS}\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}\right\rVert_{{\mathbf{G}}^{-1}}+\frac{\lambda}{\sqrt{1+2RS}}\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{\mathbf{G}}^{-1}}$
$\leq\sqrt{1+2RS}\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}\right\rVert_{{\mathbf{G}}^{-1}}+2S\sqrt{\lambda}$
$\leq 3RS\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}+2S\sqrt{\lambda}$ (Using (12) and assuming $RS\geq 1$)
Now by the optimality condition on $\widehat{\theta}$ we have $\sum_{j=1}^{s}\mu\left(\langle x_{j},\widehat{\theta}\rangle\right)x_{j}=\sum_{j=1}^{s}r_{j}x_{j}$ (see equation (3) Filippi et al. [2010]). Hence, we write
$\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}=\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}$ (13)
@@
This result is derived directly from Lemma A.2 and the definition of $\gamma(\lambda)$ (see 9). By applying the lemma, we obtain
@@
Considering the definition of $\kappa$, we have $\dot{\mu}\left(\langle\mathbf{x},\theta^{*}\rangle\right)\geq\frac{1}{\kappa}$. This implies that ${\mathbf{H}}^{*}\succeq\frac{1}{\kappa}\mathbf{V}$  [^11] For logistic rewards, we note that instead of using the worst case bound of $\kappa$, one can make use of an upper bound on $S\coloneqq\left\lVert\theta\right\rVert$ as done in Mason et al. [2022] That is, we lower bound $\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)\geq\dot{\mu}\left({S\left\lVert x\right\rVert}\right)$ and use ${\mathbf{H}}^{*}\succeq\sum_{t}\dot{\mu}\left({S\left\lVert x_{t}\right\rVert}\right)x_{t}x_{t}^{\intercal}$. which in turn leads to the inequality $\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}\leq\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$. ∎
@@ A.3 Preliminary Lemmas
This result is a direct consequence of Lemma A.2 and the definition of $\gamma(\lambda)$ (see 9). According to the lemma, we have
@@
$\leq 4\sqrt{\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}^{2}\right]\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{\mathbf{x}\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}^{2}\right]}$ (via Jensen’s inequality)
@@
$\gamma(\lambda)\leq\gamma\quad\text{ where }\gamma=30RS\sqrt{d\log{T}}.$
@@ A.5 Optimal Design Guarantees
${\mathbf{W}}_{k}$ $\coloneqq\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\operatorname*{\mathbb{E}}_{x\sim\pi_{k}({\mathcal{X}})}\left[\widetilde{x}\widetilde{x}^{\mathsf{T}}|{\mathcal{X}}\right]\right]$
@@
Lemma A.15 (Matrix Chernoff Tropp et al. [2015], Ruan et al. [2021]).
@@
${\mathbf{H}}_{k}\succeq\frac{\tau_{k}}{8}\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\operatorname*{\mathbb{E}}_{\widetilde{x}\sim\pi_{k}({\mathcal{X}})}\widetilde{x}\widetilde{x}^{\mathsf{T}}.$
@@
We note that the analysis of Ruan et al. [2021] gives an optimal guarantee (in expectation) on $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$, but not on $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}^{2}$. We obtain such a bound here and use it in the analysis.
@@
$\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\leq\frac{8}{\tau_{1}}\left\lVert x\right\rVert_{{\mathbf{W}}_{G}^{-1}}$
$\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k}^{-1}}\leq\frac{8}{\tau_{1}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{W}}_{k}^{-1}}$
We obtain the first two inequalities, (18) and (19), by a direct use of Corollary A.16. For (20) we note that for every phase we have at least $O(\sqrt{T})$ samples for learning the Distributional optimal design policy (for any M). Since, $T\geq d^{32}\log{2T}^{2}$ the event stated in Lemma A.13 holds with probability greater than $1-\frac{1}{T^{2}}$. ∎
@@ Appendix B Regret Analysis of RS-GLinCB
$\gamma\coloneqq\textbf{c}RS\sqrt{d\log{\frac{T}{\delta}}}$
@@ B.1 Confidence Sets for Switching Criterion I rounds
At any round $t$, let $\widehat{\theta}_{o}$ be the maximum likelihood estimate calculated using set of rewards observed in the rounds ${\mathcal{T}}_{o}$. With probability at least $1-\delta$ we have
@@
Let us define the matrix ${\mathbf{G}}_{w}=\sum_{s\in{\mathcal{T}}_{o}}\alpha(x,\theta^{*},\widehat{\theta}_{o})x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$. First, we note that by self-concordance property of $\mu$ (Lemma C.2), ${\mathbf{G}}_{w}\succeq\frac{1}{1+2RS}{\mathbf{H}}^{*}_{w}$. Hence,
$\lVert\widehat{\theta}_{o}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}$ $\leq\sqrt{1+2RS}\lVert\widehat{\theta}_{o}-\theta^{*}\rVert_{{\mathbf{G}}_{w}}$
$=\sqrt{1+2RS}\left\lVert{\mathbf{G}}_{w}\left(\widehat{\theta}_{o}-\theta^{*}\right)\right\rVert_{{\mathbf{G}}_{w}^{-1}}$
$=\sqrt{1+2RS}\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(\langle\widehat{\theta}_{o},x_{s}\rangle-\langle\theta^{*},x_{s}\rangle\right)\alpha(x_{s},\widehat{\theta}_{o},\theta^{*})x_{s}+\lambda\widehat{\theta}_{o}-\lambda_{t}\theta^{*}\right\rVert_{{\mathbf{G}}_{w}^{-1}}$
$=\sqrt{1+2RS}\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(\mu\left(\langle\widehat{\theta}_{o},x_{s}\rangle\right)-\mu\left(\langle\theta^{*},x_{s}\rangle\right)\right)x_{s}+\lambda\widehat{\theta}_{o}-\lambda_{t}\theta^{*}\right\rVert_{{\mathbf{G}}_{w}^{-1}}$ (Taylor’s theorem)
$\leq(1+2RS)\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(\mu\left(\langle\widehat{\theta}_{o},x_{s}\rangle\right)-\mu\left(\langle\theta^{*},x_{s}\rangle\right)\right)x_{s}+\lambda\widehat{\theta}_{o}-\lambda\theta^{*}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}$ (${\mathbf{G}}_{w}\succeq\frac{1}{1+2RS}{\mathbf{H}}^{*}_{w}$)
Since $\widehat{\theta}_{o}$ is the maximum likelihood estimate, by optimality condition, we have the following relation: $\sum_{s\in{\mathcal{T}}_{o}}\mu\left(\langle x_{s},\widehat{\theta}_{o}\rangle\right)x_{s}+\lambda\widehat{\theta}_{o}=\sum_{s\in{\mathcal{T}}_{o}}r_{s}x_{s}$. Substituting this above, we get
$\lVert\widehat{\theta}_{o}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}$ $\leq(1+2RS)\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(r_{s}-\mu\left(\langle\theta^{*},x_{s}\rangle\right)\right)x_{s}-\lambda\theta^{*}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}$
$\leq(1+2RS)\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(r_{s}-\mu\left(\langle\theta^{*},x_{s}\rangle\right)\right)x_{s}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}+\lambda\left\lVert\theta^{*}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}$
$\leq(1+2RS)\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\eta_{s}x_{s}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}+S\sqrt{\lambda}.$ (${\mathbf{H}}^{*}_{w}\succeq\lambda_{t}{\mathbf{I}}$, $\left\lVert\theta^{*}\right\rVert_{2}\leq S$)
where $\eta_{s}\coloneqq\left(r_{s}-\mu\left(\langle\theta^{*},x_{s}\rangle\right)\right)$.
We will now apply Lemma B.1 $\eta_{s}$ scaled by $R$. First note that $\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\eta_{s}x_{s}\right\rVert_{({\mathbf{H}}^{*}_{w})^{-1}}=\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\frac{\eta_{s}}{R}x_{s}\right\rVert_{(R^{2}{\mathbf{H}}^{*}_{w})^{-1}}$ which, in turn ensures that the noise variable is upper bounded by $1$. Applying B.1 we get
$\lVert\widehat{\theta}_{o}-\theta^{*}$ $\rVert_{{\mathbf{H}}^{*}_{w}}\leq S\sqrt{R^{2}\lambda}$
$+(1+2RS)\left(\frac{\sqrt{R^{2}\lambda}}{2}+\frac{2d}{\sqrt{R^{2}\lambda}}\log\left(1+\frac{\tau_{o}}{R^{2}\lambda d}\right)+\frac{2}{\sqrt{R^{2}\lambda}}\log\left(\frac{1}{\delta}\right)+\frac{2}{\sqrt{R^{2}\lambda}}d\log(2)\right)$
Simplifying constants and setting $\sqrt{R^{2}\lambda}=\textbf{c}\sqrt{d\log(T)+\log(1/\delta)}$, we have $\left\lVert\widehat{\theta}_{o}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}_{w}}\leq\textbf{c}RS\sqrt{d\log(T/\delta)+\log(1/\delta)}\leq\textbf{c}RS\sqrt{d\log{(T/\delta)}}$. ∎

### v5

@@ A.1 Additional Notation
$\gamma(\lambda)=24RS\left(\sqrt{\log{\left(T\right)}+d}+\frac{R\left(\log{\left(T\right)}+d\right)}{\sqrt{\lambda}}\right)+S\sqrt{(1+2RS)\lambda}.$ (10)
@@ A.2 Concentration Inequalities and Confidence Intervals
$\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{{\mathbf{H}}^{*}}}\leq 24RS\left(\sqrt{\log{\left(T\right)}+d}+\frac{R\left(\log{\left(T\right)}+d\right)}{\sqrt{\lambda}}\right)+S\sqrt{(1+2RS)\lambda}$ (12)
@@
$\alpha(x,\theta^{*},\widehat{\theta})$ $\coloneqq\int_{v=0}^{1}\dot{\mu}\left(\langle x,\theta^{*}\rangle+v\langle x,~{}\rangle{\left(\widehat{\theta}-\theta^{*}\right)}\right)dv$
${\mathbf{G}}$ $\coloneqq\sum_{j=1}^{s}\alpha(x,\theta^{*},\widehat{\theta})x_{j}x_{j}^{\mathsf{T}}+\lambda{\mathbf{I}}$
@@
$\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{{\mathbf{H}}^{*}}}$ $\leq\sqrt{1+2RS}\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{\mathbf{G}}}$
@@
$=\sqrt{1+2RS}\left\lVert\sum_{j=1}^{s}\left(\langle\theta^{*},x_{j}\rangle-\langle\widehat{\theta},x_{j}\rangle\right)\alpha(x,\theta^{*},\widehat{\theta})x_{j}+{\lambda}\left(\theta^{*}-\widehat{\theta}\right)\right\rVert_{{\mathbf{G}}^{-1}}$
$\leq\sqrt{1+2RS}\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}+{\lambda}\left(\theta^{*}-\widehat{\theta}\right)\right\rVert_{{\mathbf{G}}^{-1}}$ (14)
$\leq\sqrt{1+2RS}\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}-{\lambda}\widehat{\theta}\right\rVert_{{\mathbf{G}}^{-1}}+\sqrt{1+2RS}\lambda\left\lVert\theta^{*}\right\rVert_{{\mathbf{G}}^{-1}}$ (15)
$\leq(1+2RS)\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}-{\lambda}\widehat{\theta}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}+S\sqrt{(1+2RS)\lambda}$ (16)
Now by the optimality condition on $\widehat{\theta}$ we have $\sum_{j=1}^{s}\mu\left(\langle x_{j},\widehat{\theta}\rangle\right)x_{j}+\lambda~{}\widehat{\theta}=\sum_{j=1}^{s}r_{j}x_{j}$. Hence, we write
$\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}-\lambda~{}\widehat{\theta}~{}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}$ $=\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}~{}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}$ (17)
@@
This result is derived directly from Lemma A.2 and the definition of $\gamma(\lambda)$ (see 10). By applying the lemma, we obtain
@@
Considering the definition of $\kappa$, we have $\dot{\mu}\left(\langle\mathbf{x},\theta^{*}\rangle\right)\geq\frac{1}{\kappa}$. This implies that ${\mathbf{H}}^{*}\succeq\frac{1}{\kappa}\mathbf{V}$  [^11] For logistic rewards, we note that instead of using the worst case bound of $\kappa$, one can make use of an upper bound on $S\coloneqq\left\lVert\theta\right\rVert$ as done in [20] That is, we lower bound $\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)\geq\dot{\mu}\left({S\left\lVert x\right\rVert}\right)$ and use ${\mathbf{H}}^{*}\succeq\sum_{t}\dot{\mu}\left({S\left\lVert x_{t}\right\rVert}\right)x_{t}x_{t}^{\intercal}$. which in turn leads to the inequality $\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}\leq\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$. ∎
@@ A.3 Preliminary Lemmas
This result is a direct consequence of Lemma A.2 and the definition of $\gamma(\lambda)$ (see 10). According to the lemma, we have
@@
$\leq 4\sqrt{\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}^{2}\right]\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{\mathbf{x}\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}^{2}\right]}$ (via Cauchy–Schwarz)
@@
$\gamma(\lambda)\leq\gamma\quad\text{ where }\gamma=30(RS)^{3/2}\sqrt{d\log{T}}.$
@@ A.5 Optimal Design Guarantees
${\mathbf{W}}_{k}$ $\coloneqq\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\operatorname*{\mathbb{E}}_{\widetilde{x}\sim\pi_{k}(\widetilde{{\mathcal{X}}})}\left[\widetilde{x}\widetilde{x}^{\mathsf{T}}|{\mathcal{X}}\right]\right]$
@@
Lemma A.15 (Matrix Chernoff [27, 23]).
@@
${\mathbf{H}}_{k}\succeq\frac{\tau_{k}}{8}\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\operatorname*{\mathbb{E}}_{\widetilde{x}\sim\pi_{k}(\widetilde{{\mathcal{X}}})}\widetilde{x}\widetilde{x}^{\mathsf{T}}.$
@@
We note that the analysis of [23] gives an optimal guarantee (in expectation) on $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$, but not on $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}^{2}$. We obtain such a bound here and use it in the analysis.
@@
$\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\leq\frac{8}{\sqrt{\tau_{1}}}\left\lVert x\right\rVert_{{\mathbf{W}}_{G}^{-1}}$
$\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k}^{-1}}\leq\frac{8}{\sqrt{\tau_{k}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{W}}_{k}^{-1}}$
We obtain the first two inequalities, (22) and (23), by a direct use of Lemma A.12 and Lemma A.14. For (24) we note that for every phase we have at least $O(\sqrt{T})$ samples for learning the Distributional optimal design policy (for any M). Since, $T\geq d^{32}\log{2T}^{2}$ the event stated in Lemma A.13 holds with probability greater than $1-\frac{1}{T^{2}}$. ∎
@@ Appendix B Regret Analysis of RS-GLinCB
$\gamma\coloneqq\textbf{c}R^{2}S^{\nicefrac{{3}}{{2}}}\sqrt{d\log{\left(\frac{ST}{\delta}\right)}}$
@@ B.1 Confidence Sets for Switching Criterion I rounds
Let us define $g_{o}(\theta)=\sum_{s\in{\mathcal{T}}_{o}}\mu\left(\langle x_{s},\theta\rangle\right)x_{s}+\lambda\theta$ and ${\mathbf{H}}_{w}(\theta)=\sum_{s\in{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$.
@@
At any round $t$, let $\widetilde{\theta}_{o}$ be the unconstrained maximum likelihood estimate calculated using set of rewards observed in the rounds ${\mathcal{T}}_{o}$ (Algorithm 2 Line 7). With probability at least $1-\delta$ we have,
$\left\lVert g(\widetilde{\theta}_{o})-g(\theta^{*})\right\rVert_{{\mathbf{H}}_{w}^{*-1}}\leq\textbf{c}RS\sqrt{d\log(ST/\delta)}~{}.$
Proof.
First note that first order optimality condition of the MLE, we have $g(\widetilde{\theta}_{o})=\sum_{s\in{\mathcal{T}}_{o}}r_{s}x_{s}$. Therefore,
$\left\lVert g(\widetilde{\theta}_{o})-g(\theta^{*})\right\rVert_{{\mathbf{H}}_{w}^{*-1}}$ $=\left\lVert\sum_{s\in{\mathcal{T}}_{o}}r_{s}x_{s}-\sum_{s\in{\mathcal{T}}_{o}}\mu\left(\langle x_{s},\theta^{*}\rangle\right)x_{s}-\lambda\theta^{*}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}$
$\leq\left\lVert\sum_{s\in{\mathcal{T}}_{o}}(r_{s}-\mu\left(\langle x_{s},\theta^{*}\rangle\right))x_{s}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}+\left\lVert\lambda\theta^{*}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}$
$\leq\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\eta^{\prime}_{s}x_{s}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}+S\sqrt{\lambda}$ (${\mathbf{H}}_{w}^{*}\succeq\lambda{\mathbf{I}}$ and $\left\lVert\theta^{*}\right\rVert\leq S$; $\eta^{\prime}_{s}\coloneqq r_{s}-\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)$)
$=\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\frac{\eta^{\prime}_{s}}{R}x_{s}\right\rVert_{(\frac{1}{R^{2}}{\mathbf{H}}^{*}_{w})^{-1}}+S\sqrt{\lambda}$
$\leq\left(\frac{\sqrt{\nicefrac{{\lambda}}{{R^{2}}}}}{2}+\frac{2d}{\sqrt{\nicefrac{{\lambda}}{{R^{2}}}}}\log\left(1+\frac{\tau_{o}}{d\left(\nicefrac{{\lambda}}{{R^{2}}}\right)}\right)+\frac{2}{\sqrt{\nicefrac{{\lambda}}{{R^{2}}}}}\log\left(\frac{1}{\delta}\right)+\frac{2}{\sqrt{\nicefrac{{\lambda}}{{R^{2}}}}}d\log(2)\right)+S\sqrt{\lambda}$ (Lemma B.1 with $\eta_{s}=\frac{\eta^{\prime}_{s}}{R}$)
Simplifying constants and setting $\sqrt{\frac{\lambda}{R^{2}}}=\textbf{c}\sqrt{d\log(T/\delta)/S}$, we have,
$\left\lVert g(\widetilde{\theta}_{o})-g(\theta^{*})\right\rVert_{{\mathbf{H}}^{*-1}}$ $\leq\textbf{c}R\sqrt{dS\log(ST/\delta)}~{}.$
Hereon, we will fix $\lambda=\textbf{c}R\sqrt{dS\log(ST/\delta)}$ as shown in the proof above. Recall that $\widehat{\theta}_{o}$ is obtained from $\widetilde{\theta}_{o}$ via the following projection:
$\widehat{\theta}_{o}=\operatorname*{arg\,min}_{\left\lVert\theta\right\rVert\leq S}\left\lVert g_{o}(\theta)-g_{o}(\widetilde{\theta}_{o})\right\rVert_{{\mathbf{H}}_{w}(\theta)^{-1}}~{}.$
Lemma B.3.
At any round $t$, let $\widehat{\theta}_{o}$ be the projection of the maximum likelihood estimate $\widetilde{\theta}_{o}$ calculated using set of rewards observed in the rounds ${\mathcal{T}}_{o}$ (Algorithm 2, Lines 7-8). With probability at least $1-\delta$ we have
@@
Observe that $\lvert\langle x_{t},\theta^{*}-\widehat{\theta}_{o}\rangle\rvert\leq 2S$ for all $t\in{\mathcal{T}}_{o}$. Hence, by Lemma E.1 with $c=2RS$,
$\lVert\widehat{\theta}_{o}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}$ $\leq 2(1+2RS)\textbf{c}R\sqrt{dS\log(ST/\delta)}+2R\sqrt{S(1+2RS)}\leq\textbf{c}R^{2}S^{\nicefrac{{3}}{{2}}}\sqrt{d\log(ST/\delta)}~{}.$